OpenStudy (anonymous):

I don't understand the answer to 1H-1b)--how/why was the assumption made?

OpenStudy (phi):

Their explanation is a bit muddled. They are saying that lambda has the value found from part a): the half-life (given the name lambda) is \( \lambda= \frac{\ln 2}{k} \) For part 2) we are given \[ y_1= y_0\ e^{-k t_1} \] now find y at time \(t_1 + \lambda\): \[y_2 = y_0\ e^{-k (t_1+ \lambda)} \\ = y_0\ e^{-k t_1}e^{-k \lambda} \\ = y_1\ e^{-k \cdot \frac{\ln 2}{k} } \\ = y_1 \cdot \frac{1}{e^{\ln 2}} \\ y_2= \frac{y_1}{2} \]